Breakeven cash inflows. The One Ring Company​, a leading producer of fine cast silver​ jewelry, is considering the purchase of new casting equipment that will allow it to expand its product line. The​ up-front cost of the equipment is $744,000. The company expects that the equipment will produce steady income throughout its 11​-year life.  

a. If One Ring requires a 10 % return on its​ investment, what minimum yearly cash inflow will be necessary for the company to go forward with this​ project?

b. How would the minimum yearly cash inflow change if the company required a 13% return on its​ investment?
Answer #1

a) This part of the question is based on the concept of Present value of Time value of money concept which states that Present value = Annual Inflow * PVF ( r%, t)

Applying the above formula:

Annual inflow = Present value / PVF (r%,t) where Present value = 744000 , r = 10% and t = 11 years

Annual inflow = 744,000 / 6.495061

Annual inflow = $ 114,548.58

Minimum yearly cash flow that is necessary to work further is $ 114,548.58

b) The minimum yearly cash flow if the rate of required return is 13% is :

Annual inflow = 744,000 / PVF(13%,11)

Annual inflow = 744000/ 5.686941

Annual inflow = $ 130,826.04

Minimum yearly cash flow that is necessary to work further is $ 130,826.04 if the rate of return is 13%